A rectangular field is four times as long as it is wide. If the perimeter of the field is 1000 feet, what are the dimensions of the field?
A) Write an equation you can use to answer the given question. Let w be the width of the field.

Answer by ewatrrr(24785)

 
Hi,

P = 2L + 2w = 1000ft
0r L + w = 500ft
Question States*** L = 4w
4w + w = 500ft
5w = 500ft
w = 100ft and L is 400ft
dimensions of field are 400ft by 100ft